A piano that weighs 5000N is pushed up a ramp inclined 12º to horizontal. If the applied
force is 1500N and frictional force is 50N,calculate the net force on that piano.
1
Expert's answer
2020-04-06T09:15:40-0400
"F_{net}=F-W\sin{12}-F_f"

"F_{net}=1500-5000\sin{12}-50=410\ N"